OpenStudy (anonymous):

What is the equation for the vertical asymptotes of the function f(x)=-3tan(0.5x)?

OpenStudy (anonymous):

I have one equation being y=2npi+pi is this right?

OpenStudy (psymon):

Well, do you know the normal period of the tangent function?

OpenStudy (anonymous):

pi

OpenStudy (psymon):

Right. And do you know where the tangent asymptotes normally are?

OpenStudy (anonymous):

y=(npi/2)

OpenStudy (psymon):

Alright, well when you have a function atan(bx + c), the new period of the function is pi/b. So in this case, your b is .5, making your new period 2 pi. Now as for where your asymptotes start from. We have two asymptoes, one at -pi/2 and the other at pi/2. Now we make an inequality with our tangent angle in between these two asymptotes like this: \[-\frac{ \pi }{ 2 }\le.5x \le \frac{ \pi }{ 2 }\] Now we solve for x and that will give us two asymptotes from which we can add our period on to to get all of our asymptotes.

OpenStudy (anonymous):

y=npi? is the asympyotes?

OpenStudy (psymon):

Your asymptotes start at -pi and then add 2npi. because 2pi is our new period.

OpenStudy (anonymous):

Oh!! I get it! Thanks!

OpenStudy (psymon):

Awesome xD Np :3